Calculate the concentrations of all species in a 0.270-M solution of H2X if K1 = 6.31E-06 and K2 = 1.74E-08

[H2X] =

[H3O1+] =

[HX1-] =

[X2-] =
Answer #1

                  H2O + H2X <====> H3O+ + HX-                             Ka1 = 6.31 x 10-06

Intial Con               0.270              10-7     0

Change Con             -x                  +x       +x

Eq Con                  0.27-x          10-7+x    +x

Ka1 = (10-7+x)(x) / ( 0.27-x ) => 6.31 x 10-06     ( x >> 10-7 or simply x = [ Ka1*C]1/2 )

x = 1.31 x 10-3 M

x = [H3O+] = [ HX-] = 1.31 x 10-3 M

[H2X ] = 0.270 - 0.00131 => 0.2687 M

                    H2O + HX-             <====>    H3O+ +            X2-                             Ka2 = 1.74 x 10-08

Intial Con               1.31 x 10-3                            1.31 x 10-3              0

Change Con             -x                               +x                       +x

Eq Con                  1.31 x 10-3 -x           1.31 x 10-3 +x            +x

Ka2 = (1.31 x 10-3 +x)(x) / ( 1.31 x 10-3 -x  ) => 1.74 x 10-08

x = 1.74 x 10-08 M

x = [ X2-] = 1.74 x 10-08 M

( Note in diprotic acid [ X2-] =Ka2 )

[H2X] = 0.2687 M

[H3O1+] = 1.31 x 10-3 M

[HX1-] = 1.31 x 10-3 M

[X2-] = 1.74 x 10-08 M